The supply and demand for broccoli are described by the following equations:

Supply: Qss = 4P - 80

Demand: Qdd = 100 - 2P

Q is in bushels, and P is in dollars per bushel.

a. Graph the supply curve and the demand curve. What is the equilibrium price and quantity?
Answer #1

The equailibrium quantity and price is determined when the quantity supplied is equal to quantity demanded.

We can see QSS= 4P-80, this is in form y=mx+c slope line equation where m is the slope and here it is positive(4P).

Qdd= 100-2P, this is in form y=mx+c slope line equation. and here m is negative -2P.

So, equailibrium will be 4P-80=100-2P,

6P=180; P=30.

Keeping P in these equation of Qss and Qdd. Qss=40; Qdd=40.

ส H -- - -- - - <